# Short Paper: Atomic Execution is Not Enough for Arbitrage Profit Extraction in Shared Sequencers

### v2

@@ 1 Introduction
Decentralized Finance (DeFi)has been essential to the growth of the Ethereum ecosystem, attracting many users and successful applications. Recently, it has expanded to Layer-2 (L2)scaling solutions like rollups, where trading volumes are rising, with some rollups now experiencing more daily activity than Ethereum itself [4]. With this growth in adoption comes more opportunities for Maximal-Extractable Value (MEV) — a collection of techniques for extracting value from transaction inclusion and reordering [3]. The most prevalent form of MEVon rollups is arbitrage, in which arbitrageurs exploit price differences between centralized exchanges and/or Decentralized Exchanges (DEXs) [10].
A relevant consideration for MEVin rollups is sequencer design. The sequencer is the operator responsible for receiving and scheduling user transactions for processing, and most rollups currently use an independent centralized sequencer [6]. Recent proposals have introduced an alternative — shared sequencers. Shared sequencing schemes propose to process transactions for multiple rollups together, allowing for better composability between rollups. Despite being a recent topic, Astria [1] already has a solution in production, while Radius [8], NodeKit [7], and Expresso Systems [9, 2] are in the test phase. However, we have not yet seen significant adoption from rollups.
@@ 2.3 Trade Sizes
In the swap $S_{B}$, the arbitrageur sends $\Delta y_{B}$ units of token $Y$ to the pool and pays a fee of $f$. Here, we consider that the DEXis processing fees outside of the pool reserves, which means that when a trader wishes to swap a given amount of tokens and pay $\Delta y$, only part of this payment goes to the pool reserve. Concretely, $(1-f)\cdot\Delta y$ is added to the poor reserves, while $f\cdot\Delta y$ is paid to Liquidity Providers (LPs). This is the case of Uniswap V3 pools, for instance.
@@
$P_{B}^{\text{end}}$ $=\frac{y_{B}+(1-f)\Delta y_{B}}{x_{B}-\Delta x_{B}}=\frac{y_{B}+(1-f)\Delta y_{B}}{x_{B}-\frac{x_{B}(1-f)\Delta y_{B}}{y_{B}+(1-f)\Delta y_{B}}}=$
$=\frac{\left[y_{B}+(1-f)\Delta y_{B}\right]^{2}}{x_{B}\cdot y_{B}+x_{B}(1-f)\Delta y_{B}-x_{B}(1-f)\Delta y_{B}}=$
$=\frac{\left[y_{B}+(1-f)\Delta y_{B}\right]^{2}}{x_{B}\cdot y_{B}}$ (7)
@@
$(1-f)^{2}P_{A}^{\text{end}}=P_{B}^{\text{end}}\Longleftrightarrow$
$\Longleftrightarrow(1-f)^{2}\frac{x_{A}y_{A}}{\left[x_{A}+(1-f)\Delta x_{A}\right]^{2}}=\frac{\left[y_{B}+(1-f)\Delta y_{B}\right]^{2}}{x_{B}y_{B}}$
$\Longleftrightarrow\frac{(1-f)\sqrt{x_{A}y_{A}}}{x_{A}+(1-f)\Delta x_{A}}=\frac{y_{B}+(1-f)\Delta y_{B}}{\sqrt{x_{B}y_{B}}}$
$\Longleftrightarrow\frac{(1-f)\sqrt{x_{A}y_{A}}}{x_{A}+(1-f)\frac{x_{B}(1-f)\Delta y_{B}}{y_{B}+(1-f)\Delta y_{B}}}=\frac{y_{B}+(1-f)\Delta y_{B}}{\sqrt{x_{B}\cdot y_{B}}}$
$\Longleftrightarrow\frac{(1-f)\sqrt{x_{A}\cdot y_{A}}}{x_{A}\left[y_{B}+(1-f)\Delta y_{B}\right]+(1-f)x_{B}(1-f)\Delta y_{B}}=\frac{1}{\sqrt{x_{B}\cdot y_{B}}}$
$\Longleftrightarrow\frac{(1-f)\sqrt{x_{A}\cdot y_{A}}}{x_{A}\cdot y_{B}+x_{A}(1-f)\Delta y_{B}+x_{B}(1-f)^{2}\Delta y_{B}}=\frac{1}{\sqrt{x_{B}\cdot y_{B}}}$
$\Longleftrightarrow x_{A}\cdot y_{B}+x_{A}(1-f)\Delta y_{B}+x_{B}(1-f)^{2}\Delta y_{B}=(1-f)\sqrt{x_{A}\cdot y_{A}\cdot x_{B}\cdot y_{B}}$
$\Longleftrightarrow\Delta y_{B}=\frac{(1-f)\sqrt{x_{A}\cdot y_{A}\cdot x_{B}\cdot y_{B}}-x_{A}\cdot y_{B}}{(1-f)x_{A}+(1-f)^{2}x_{B}}$ (10)

### v3

@@ 1 Introduction
Decentralized Finance (DeFi)has been essential to the growth of the Ethereum ecosystem, attracting many users and successful applications. Recently, it has expanded to Layer-2 (L2)scaling solutions like rollups, where trading volumes are rising, with some rollups now experiencing more daily activity than Ethereum itself [4]. With this growth in adoption comes more opportunities for Maximal-Extractable Value (MEV) — a collection of techniques for extracting value from transaction inclusion and reordering [3]. One of the most prevalent forms of MEVon rollups is arbitrage, in which arbitrageurs exploit price differences between centralized exchanges and/or Decentralized Exchanges (DEXs) [14, 15].
A relevant consideration for MEVin rollups is sequencer design. The sequencer is the operator responsible for receiving and scheduling user transactions for processing, and most rollups currently use an independent centralized sequencer [8]. Recent proposals have introduced an alternative — shared sequencers. Shared sequencing schemes propose to process transactions for multiple rollups together, allowing for better composability between rollups. Despite being a recent topic, Astria [1] already has a solution in production, while Radius [11], NodeKit [9], and Expresso Systems [12] are in the test phase. However, we have not yet seen significant adoption from rollups.
@@ 2.3 Trade Sizes
In the swap $S_{B}$, the arbitrageur sends $\Delta y_{B}$ units of token $Y$ to the pool and pays a fee of $f$. Here, we consider that the DEXis processing fees outside of the pool reserves, which means that when a trader wishes to swap a given amount of tokens and pay $\Delta y$, only part of this payment goes to the pool reserve. Concretely, $(1-f)\cdot\Delta y$ is added to the pool reserves, while $f\cdot\Delta y$ is paid to Liquidity Providers (LPs). This is the case of Uniswap V3 pools, for instance.
@@
$P_{B}^{\text{end}}=\frac{y_{B}+(1-f)\Delta y_{B}}{x_{B}-\Delta x_{B}}=\frac{y_{B}+(1-f)\Delta y_{B}}{x_{B}-\frac{x_{B}(1-f)\Delta y_{B}}{y_{B}+(1-f)\Delta y_{B}}}=\frac{\left[y_{B}+(1-f)\Delta y_{B}\right]^{2}}{x_{B}\cdot y_{B}}$ (7)
@@
$(1-f)^{2}P_{A}^{\text{end}}=P_{B}^{\text{end}}\Longleftrightarrow\Delta y_{B}=\frac{(1-f)\sqrt{x_{A}\cdot y_{A}\cdot x_{B}\cdot y_{B}}-x_{A}\cdot y_{B}}{(1-f)x_{A}+(1-f)^{2}x_{B}}$ (10)
@@ 3 Atomicity Profit Conditions
Focusing on the failing probabilities, there is a special case we can analyze. If the two failure probabilities are equal (i.e., $f_{A}=f_{B}$), the expected profit difference is always negative, meaning that, on average, the arbitrageur should stick with the status quo without atomicity. This result comes directly from Equation 12:
$\mathbb{E}[\text{Profit}_{\text{diff}}]=$
$=\Delta x_{B}\big{[}f_{A}(P^{*}_{B}-P_{\text{ext}})+f_{A}(P_{\text{ext}}-P^{*}_{A})+f_{A}^{2}(P^{*}_{A}-P^{*}_{B})\big{]}=$
$=\Delta x_{B}\big{[}f_{A}(P^{*}_{B}-P^{*}_{A})+f_{A}^{2}(P^{*}_{A}-P^{*}_{B})\big{]}=$
$=\Delta x_{B}\cdot f_{A}\cdot(1-f_{A})\cdot(P^{*}_{B}-P^{*}_{A})<0$ (13)
4 Related Work
The seminal work from Daian et al. [3] laid the groundwork for understanding MEVby demonstrating how block producers could exploit transaction ordering on Ethereum to capture arbitrage and frontrunning profits. Since then, multiple empirical studies have examined MEVon Ethereum by uncovering common strategies and measuring their prevalence and impact [13, 10].
Recent work has extended the analysis into the Layer-2 domain. For instance, Torres et al. [14], Ha et al. [5] and Bagourd and Francois [2] present different measurements of MEV across popular rollups and other Layer-2s, revealing the most common strategies, their volume, and the corresponding profit. Complementarily, Gogol et al. [4] examines arbitrage across different rollups and identifies many untapped arbitrage opportunities resulting from the non-atomic nature of cross-rollup transactions. Finally, Öz et al. [15] systematically analyzes non-atomic cross-chain arbitrage strategies across multiple L2s. It reveals that liquidity fragmentation across heterogeneous blockchain networks creates substantial arbitrage opportunities while also highlighting the inherent challenges posed by non-atomic execution.
Beyond empirical analyses, McMenamin [7] categorizes MEV extraction methods across multiple domains and outlines proposals for shared sequencers coordinating transaction ordering across chains. In addition, Mamageishvili and Schlegel [6] models the economic incentives under different sequencing regimes. Their results suggest that while a unified sequencer could facilitate atomic cross-chain arbitrage, it may also intensify latency competition and not necessarily increase overall sequencer revenue compared to independent rollup-specific sequencers.